One mole is the amount of any given chemical substance that helps us to measure it's mass conveniently on a scale of C12 Isotope.

How many particles (atoms of carbon) are present in 12 g of C12 isotope of a naturally occuring sample of carbon.



1
Expert's answer
2022-01-28T13:00:03-0500

Mass of Carbon = 12 g .



Moles of Carbon =

mass of carbon / Atomic mass of carbon .



Moles of Carbon = 12 / 12 = 1 mol .




Number of atoms of carbon =

Moles of carbon × Avogadro number .


Number of atoms of carbon =

1 × N = 6.02 × 1023 atoms . ANSWER .